Can you explain step by step how to find R

VxAAAAAElFTkSuQmCC

Can you explain step by step how to find R
Answer #1

0 =-110+32.1/(r*(1+r))
110 =32.1/(r+r^2)
r+r^2 =32.1/110 =
r^2+r-0.291818181818182 =0
Using quadratic formula
r=(-b+(b^2-4ac)^0.5)/(2a) =(-1+(1^2-4*1*-0.291818181818182)^0.5)/(2*1) =23.61%